OpenStudy (anonymous):

the length of a rectangle is three times the width. what is the width if the area is 147 square feet

OpenStudy (anonymous):

We can say that L = 3W

OpenStudy (anonymous):

X * 3X = 147 3(X^2) = 147 X^2 = 147/3 X^2 = 49 X = sqrt(49) X = width = 7 Width 7 x Length 21 = area 147

OpenStudy (anonymous):

And we know that L * W = 147

OpenStudy (anonymous):

thank you so much